I need a solution and shown work for this problem. Help.
Solve each formula for the indicated variable.


I = E/R+r solve for R

Answer by checkley71(8403)

I=E/R+r now combine E/R & r by multiplying r by R & adding these two terms all over R or I=(E+Rr)/R now cross multiple terms to get IR=E+Rr now subtract Rr from both sides of the equation thus IR-Rr=E now factor out R or R(I-r)=E now divide both sides of the equation by (I-r) or R=E/(I-r).
hope this helps.